What is the percentage of MnCl2 in MnCl2 2H2O?
1
Expert's answer
2018-03-19T08:44:55-0400
Answer
M(MnCl 2 ) = 55 + 35.5x2 = 126 g/mol
M(MnCl 2 x2H 2 O) = (55 + 35.5x2) + (2x18) = 162 g/mol
w(MnCl 2 ) = 126/162x100% = 77.78%